OpenStudy (anonymous):

Last and Final Oneeee... :) -3=5n+12

OpenStudy (anonymous):

ok your first step is?

OpenStudy (anonymous):

+-3 on both sides

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

try again

OpenStudy (anonymous):

Umm well we can do that but the best thing to do is to keep the n on the right side. That means that everything on the right hand side must go

OpenStudy (anonymous):

except for the n

OpenStudy (anonymous):

yeah umm -12 on both sides

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

good job now what's the next step

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

and then you have your answer

OpenStudy (anonymous):

-3=5n+12 -3-12=5n+12-12 -3-12=5n -15=5n

OpenStudy (anonymous):

divide both sides by 5

OpenStudy (anonymous):

yess -5. lol that was a easy one

OpenStudy (anonymous):

you should end up with -3=n

OpenStudy (anonymous):

ooops i mean -3

OpenStudy (anonymous):

-15=5n -15/5=5n/5 -15/5=n -3=n